Compute the minimum distance that a truck slides on a horizontal asphalt road if it is traveling at 20 m/s? The coefficient to sliding friction between asphalt and rubber tire is at 0.50. The weight of the truck is 8000 kg


1
Expert's answer
2022-04-29T12:30:38-0400

The work-energy theorem says

"\frac{mv^2}{2}=Fd=\mu mg d"

Hence, the distance traveled by a truck to the rest

"d=\frac{v^2}{2\mu g}=\frac{20^2}{2*0.5*9.8}=40.8\:\rm m"